What is the difference of the values of the two variables in this system of equations?
y = 2x + 1
x + 3y = 10

Answer by ikleyn(52835)

.

Substitute  y = 2x+1  from the first equation to the second one


    x + 3*(2x+1) = 10

    
    x + 6x + 3 = 10

      7x       = 10-3 = 7

       x              = 7/7 = 1.


Then  from the first equation  y = 2*1 + 1 = 3.


The difference  x - y  is  1 - 3 = -2.

The difference  y - x  is   3 - 1 = 2.

Solved.